# Efficient Visual Transformer by Learnable Token Merging

### v1

@@ Abstract
Self-attention and transformers have been widely used in deep learning. Recent efforts have been devoted to incorporating transformer blocks into different neural architectures, including those with convolutions, leading to various visual transformers for computer vision tasks. In this paper, we propose a novel and compact transformer block, Transformer with Learnable Token Merging (LTM), or LTM-Transformer. LTM-Transformer performs token merging in a learnable scheme. LTM-Transformer is compatible with many popular and compact transformer networks, and it reduces the FLOPs and the inference time of the visual transformers while maintaining or even improving the prediction accuracy. In the experiments, we replace all the transformer blocks in popular visual transformers, including MobileViT, EfficientViT, ViT-S/16, and Swin-T, with LTM-Transformer blocks, leading to LTM-Transformer networks with different backbones. The LTM-Transformer is motivated by reduction of Information Bottleneck, and a novel and separable variational upper bound for the IB loss is derived. The architecture of mask module in our LTM blocks which generate the token merging mask is designed to reduce the derived upper bound for the IB loss. Extensive results on computer vision tasks evidence that LTM-Transformer renders compact and efficient visual transformers with comparable or much better prediction accuracy than the original visual transformers. The code of the LTM-Transformer is available at https://github.com/Statistical-Deep-Learning/LTM.
@@ 1 Introduction
Building upon the success of Transformer in natural language processing  [1],visual transformers have demonstrated remarkable performance across a wide range of tasks [2, 3, 4, 5, 6, 7]. However, the achievements of visual transformers are accompanied with heavy computational costs [3, 8], making their deployment impractical under resource-limited scenarios. The aforementioned limitations have spurred recent research endeavors aimed at developing efficient visual transformers. In this paper, we study the problem of accelerating visual transformers by token merging.
@@
Motivation. Due to the fact that the FLOPs of a visual transformer largely depend on the number of tokens in all the transformer blocks, the FLOPs of a visual transformer can be significantly reduced by reducing the number of tokens in all the transformer blocks. Our goal is to merge the output tokens of all the transformer blocks into less tokens without largely sacrificing the prediction accuracy of the original visual transformer. However, directly merging the output tokens, even by carefully designed methods [12, 13, 14], would adversely affect the performance of the model. In this paper, we propose to maintain a compelling prediction accuracy of a visual transformer with token merging by an informative token merging process. In our LTM-Transformer block, the original attention output tokens of a transformer block are merged into less target tokens, and every target token is an informative weighted average of the original output tokens. All the target tokens, or merged tokens are the final attention output tokens for the LTM-Transformer block, which are fed to an MLP to produce the output of the LTM-Transformer block as illustrated by Figure 1.
@@
Such an idea of informative token merging can also be viewed from the perspective of Information Bottleneck (IB). Let $Z$ be the original attention output tokens, which are merged into the merged tokens denoted by $\tilde{X}$, and let $Y$ be the ground truth training labels for a classification task. $\tilde{X}$ has less tokens than $Z$. The principle of IB is to maximize the mutual information between $\tilde{X}$ and $Y$ while minimizing the mutual information between $\tilde{X}$ and $X$. That is, IB encourages the network to learn the merged tokens more correlated with the class labels while reducing their correlation with the input. Extensive empirical and theoretical works have evidenced that models respecting the IB principle enjoy compelling generalization. With the informative token merging process in LTM-Transformer, the merged tokens $\tilde{X}$ are the informative aggregation of the original attention output tokens $Z$, so $\tilde{X}$ are less correlated with the training images and in this manner the IB principle is better adhered. This is reflected in Table V in Section 4.4, where a model for ablation study with token merging, ToMe, enjoys less IB loss than the vanilla transformer, MobileViT-S. This observation indicates that the IB principle is better respected by the token merging process in ToMe. In order to further decrease the IB loss, we propose an Information Bottleneck (IB) inspired token merging process, where a LTM-Transformer block generates an informative token merging task which reduces the IB loss for visual transformers. For example, our model termed “LTM-MobileViT-S” in Table V is the visual transformer with the IB loss reduced by replacing all the transformer blocks in MobileViT-S with LTM-Transformer blocks so that more informative merged tokens are generated by the proposed informative token merging process. While ToMe hurts the prediction accuracy compared to the vanilla model, our LTM-Transformer enjoys even higher top-1 accuracy than the vanilla MobileViT-S, and we have the same observations for MobileViT-XS and EfficientViT.
1.1 Contributions
The contributions of this paper are presented as follows.
@@ 1.1 Contributions
First, we present a novel and compact transformer block termed Transformer with Learnable Token Merging, or LTM-Transformer. Our LTM-Transformer block generates an informative token merging mask which reduces the IB loss. The LTM-Transformer blocks can be used to replace all the transformer blocks in many popular visual transformers, rendering compact visual transformers with competitive performance. The effectiveness of LTM-Transformer is evidenced by replacing all the transformer blocks in popular visual transformers which are already compact, including MobileViT [17], EfficientViT [7], ViT-S/16 [3], and Swin-T [4], with LTM-Transformer blocks, for image classification, object detection and instance segmentation tasks.
@@
Second, we propose an informative token merging process for visual transformers which can reduces the IB loss. As a first step, we derive a novel and separable variational upper bound for the IB loss associated with token merging, which is $I(\tilde{X}(G))-I(\tilde{X}(G),Y)$ where $I(\cdot,\cdot)$ denotes mutual information and $G$ is the token merging mask in LTM-Transformer. We then view a transformer with multiple LTM-Transformer blocks as an iterative process for reduction of the IB loss by gradient descent, and every LTM-Transformer block performs one-step gradient descent on the variational upper bound for the IB loss. Inspired by this understanding, the token merging mask at the current layer is generated from the token merging mask at the previous layer and the input tokens at the current layer by a learnable mask module, following the formula of gradient descent as in (3.2) in Section 3.2. As a result, such informative token merging process generates in a network with LTM-Transformer blocks enjoys reduced IB loss, which is evidenced in our ablation study in Section 4.4. Due to the separability of the variational upper bound for the IB loss, a neural network with LTM-Transformer blocks can be trained in an end-to-end manner with standard SGD. We remark that as shown in Table V, a baseline token merging method, ToMe, can already reduce the IB loss. By replacing all the transformer blocks with our LTM-Transformer blocks, the network with LTM-Transformer enjoys smaller IB loss, higher classification accuracy and less FLOPs. Importantly, extensive experiment results on various computer vision tasks demonstrate the compelling performance of LTM-Transformer networks compared to the competing baselines.
It is worthwhile to mention that our LTM-Transformers are trained from scratch, while most existing token merging methods [14, 12, 13] perform token merging on a trained transformer without additional training [14] or with a light training process [12, 13]. The training time of LTM-Transformers with comparison to competing token merging methods is presented in Table VI of Section 5.1.
@@
This paper is organized as follows. The related works in efficient visual transformers and compression of visual transformers by pruning or token merging are discussed in Section 2. The formulation of LTM-Transformer is detailed in Section 3. The effectiveness of LTM-Transformer is demonstrated in Section 4 for image classification, object detection and instance segmentation tasks, by replacing all the transformer blocks of various popular visual transformers, including MobileViT [17], EfficientViT [7], ViT-S/16 [3], and Swin-T [4], with LTM-Transformer blocks. Our ablation studies are presented in Section 5 and Section 4.4, and the proofs of theoretical results are in Section 6. We conclude the paper in Section 7. Throughout this paper we use $[n]$ to denote the natural numbers between $1$ and $n$ inclusively.
@@ 2.1 Efficient Visual Transformers
Visual transformer models have recently achieved superior performance on a variety of computer vision applications [18, 4, 19, 5, 6, 20]. However, visual transformers often encounter high computational demands due to the quadratic complexity of the point-wise attention and numerous Multi-Layer Perceptron (MLP) layers. To mitigate the challenges of high computational costs, various strategies have been developed [5, 2], primarily aimed at refining the network architectures and incorporating sparse mechanisms for efficient computation. These include the integration of convolutions into transformer networks [17, 7, 21], the use of knowledge distillation for training more efficient transformers [22, 23, 24], and compressing existing visual transformers with methods such as pruning [25, 26, 27]. Techniques for compressing visual transformers generally fall into three categories: (1) Channel Pruning, which targets the elimination of superfluous heads and channels within ViT blocks [25, 28, 29]; (2) Block Pruning, which involves removing redundant transformer blocks [30, 26]; (3) Token Pruning and Token Merging, which prune less important tokens and merge similar tokens in the input of transformer blocks [15, 27, 14, 31].
In this paper, we focus on learning to merge tokens guided by the information bottleneck theory of deep learning and primarily review existing works on Token Pruning and Merging [31, 15, 14, 13, 12]. DynamicViT [15] observes that the prediction in visual transformers is only based on a subset of the most informative tokens and proposes a hierarchical token sparsification framework to prune redundant tokens. ToMe [14] proposes a graph-based matching algorithm that combines similar tokens in each visual transformer block of a pre-trained visual transformer. LTMP [13] learns threshold masking modules that dynamically determine which tokens to merge and prune in a unified framework similar to DynamicViT. ToFu [12] also combines token pruning and token merging. Instead of average merging similar tokens, ToFu proposes a conventional average merging module to improve the quality of merged tokens.
@@ 2.2 Related Works about Information Bottleneck
[32] provides the first in-depth analysis of conventional information bottleneck (IB) theories and deep learning to establish the connection between the nonlinearity of neural networks and the compression phase of training. Building on the theory of IB, [33] proposes a probabilistic attention module reducing mutual information between the input and the masked representation while increasing mutual information between the masked representation and the task label. Further exploring the mechanics of IB in deep learning, [34] finds that self-attention mechanisms can be interpreted as iterative steps in optimizing the IB objective, which explains the advantages of self-attention in learning robust representation. Distinct from most existing methods that implicitly incorporate the IB principle, our work adopts a direct and innovative approach. We aim to optimize a novel and separable variational upper bound of the IB loss with a learnable token merging method. The proposed LTM-Transformer lead to compelling performance on many popular visual transformer architecture with lower computation cost, benefiting from the learnable token merging mechanism guided by the IB principle.
(a) LTM-Transformer block for regular transformers, such as ViT and Swin.
(b) LTM-Transformer block for efficient transformers, such as MobileViT and EfficientViT.
@@ 3.1 Token Merging by Learnable Token Merging Masks
Given the input feature tokens $X\in\mathbb{R}^{N\times D}$ where $N$ is the number of tokens and $D$ is the token dimension, the LTM-Transformer block first applies the self-attention module on the input feature tokens by $Z=\text{ATTN}(X)$, where $\text{ATTN}(\cdot)$ is the regular QKV self-attention operation [35]. As illustrated in Figure 1, every LTM-Transformer block has an MLP as a learnable mask module which generates the token merging mask $G^{(\ell)}$ where $\ell$ is the index of the current layer or block. The LTM-Transformer block merges the $N$ tokens of $Z$ into $P$ tokens with $P<N$ by multiplying $Z$ with the token merging mask $G^{(\ell)}\in\mathbb{R}^{N\times P}$. We set $P=\lceil r\times N\rceil$, where $r\in(0,1)$ is termed the compression ratio for LTM, and a smaller $r$ renders less merged tokens after token merging. The token merging mask $G^{(\ell)}$ of the $\ell$-th transformer block is generated by the token merging mask $G^{(\ell-1)}$ of the previous layer and the feature tokens $Z$, which is motivated by reducing the IB loss and detailed in Section 3.2. The token merging mask $G^{(1)}$ for the first block is generated from the input feature tokens of the first layers by $G^{(1)}=\text{Softmax}(f(X/\tau))$, where $f(\cdot)$ represents an MLP layer followed by a Sigmoid function. $\tau\in(0,1)$ is a temperature parameter that aims to sparsify the merging mask. After obtaining the merging mask $G^{(\ell)}$, the features tokens of $Z$ are merged into $P$ tokens by $\tilde{X}(G^{(\ell)})=\left(Z^{\top}G^{(\ell)}\right)^{\top}\in\mathbb{R}^{P\times D}$, which is then passed to the following MLP layers in the transformer block.
@@
In addition to merging tokens in regular transformer blocks such as ViT [35] and Swin [4], the LTM-Transformer block can also be applied to efficient transformer blocks widely applied in efficient visual transformer architectures such as MobileViT [17] and EfficientViT [7]. Regular transformer blocks obtain the output by sequentially applying the attention operation and MLP on the input feature tokens. However, efficient transformer blocks usually contain residual connections following the design of residual connections in Convolutional Neural Networks (CNNs). That is, these blocks maintain the same shapes for the input $X$ and the self-attention output $Z$ and concatenate them to produce the output features of the current transformer block. As a result, we cannot only merge the tokens of $Z$. Instead, our LTM-Transformer block merges the tokens of both $X$ and $Z$ so that the number of merged tokens for $X$ and $Z$ have is the same. To this end, we apply the same token merging mask $G^{(\ell)}$ to merge both $X$ and $Z$. As a result, the compressed $X$ and $Z$ are of the same shape after the token merging process and they can still be concatenated, which is illustrated in Figure 1(b). In addition, transformer blocks in the efficient visual transformers are usually accompanied with convolution operations so that they need to maintain the feature tokens in a three-dimensional format $X\in\mathbb{R}^{H\times W\times D}$ as illustrated in Figure 1(b). To apply our token merging method on efficient transformer blocks, we set the number of merged tokens after token merging as $P=H^{\prime}\times W^{\prime}$, where $r$ is the compression ratio, and $H^{\prime}=\lceil H\times\sqrt{r}\rceil,W^{\prime}=\lceil W\times\sqrt{r}\rceil$. Therefore, the merged tokens can still be reshaped into three-dimensional features for later convolution operations. The analysis about the inference computation cost, or the FLOPs, of the LTM transformer block for token merging in both regular transformers and efficient transformers as illustrated in Figure 1 is detailed below.
@@
Computation Cost Analysis of LTM-Transformer for Token Marging. We hereby analyze the additional inference computation cost, or the FLOPs, of the LTM transformer block for token merging in both regular transformers and efficient transformers as illustrated in Figure 1. Let $D$ be the dimension of input tokens and $N$ be the number of tokens. The FLOPs of the token merging in an LTM transformer block in regular visual transformers is $6CDP+3C+ND^{2}+NDP$, where $6CDP+3C+ND^{2}$ is the FLOPs for calculating the merging mask and $NDP$ is the cost for applying the merging mask on the input tokens. In the LTM transformer block of efficient visual transformers, the additional FLOPs of the token merging is $6CDP+3C+ND^{2}+2NDP$, since the merging mask will be applied to both the input tokens and the merged tokens.
@@
3.2 Generating Token Merging Mask by Reducing Variational Upper Bound for the IB Loss

### v2

@@ Abstract
Self-attention and transformers have been widely used in deep learning. Recent efforts have been devoted to incorporating transformer blocks into different neural architectures, including those with convolutions, leading to various visual transformers for computer vision tasks. In this paper, we propose a novel and compact transformer block, Transformer with Learnable Token Merging (LTM), or LTM-Transformer. LTM-Transformer performs token merging in a learnable scheme. LTM-Transformer is compatible with many popular and compact transformer networks, and it reduces the FLOPs and the inference time of the visual transformers while maintaining or even improving the prediction accuracy. In the experiments, we replace all the transformer blocks in popular visual transformers, including MobileViT, EfficientViT, ViT, and Swin, with LTM-Transformer blocks, leading to LTM-Transformer networks with different backbones. The LTM-Transformer is motivated by reduction of Information Bottleneck, and a novel and separable variational upper bound for the IB loss is derived. The architecture of the mask module in our LTM blocks, which generates the token merging mask, is designed to reduce the derived upper bound for the IB loss. Extensive results on computer vision tasks evidence that LTM-Transformer renders compact and efficient visual transformers with comparable or much better prediction accuracy than the original visual transformers. The code of the LTM-Transformer is available at https://github.com/Statistical-Deep-Learning/LTM.
@@ 1 Introduction
Building upon the success of Transformer in natural language processing  [1], visual transformers have demonstrated remarkable performance across a wide range of tasks [2, 3, 4, 5, 6, 7, 8]. However, the achievements of visual transformers are accompanied with heavy computational costs [3, 9], making their deployment impractical under resource-limited scenarios. The aforementioned limitations have spurred recent research endeavors aimed at developing efficient visual transformers. In this paper, we study the problem of accelerating visual transformers by token merging.
@@
Motivation. Due to the fact that the FLOPs of a visual transformer largely depend on the number of tokens in all the transformer blocks, the FLOPs of a visual transformer can be significantly reduced by reducing the number of tokens in all the transformer blocks. As reviewed in Section 2.1, the existing token merging methods do not have a principled way of merging original tokens based on their informativeness and importance for classification tasks. Our goal is to design a principled informative token merging method to merge the output tokens of every transformer block into fewer tokens without largely sacrificing the prediction accuracy of the original visual transformer, and more informative original tokens contribute more to the merged tokens in the token merging process. However, directly merging the output tokens, even by carefully designed methods [13, 14, 15], would adversely affect the performance of the model. In this paper, we propose to maintain a compelling prediction accuracy of a visual transformer with token merging by an informative token merging process. In our LTM-Transformer block, the original attention output tokens of a transformer block are merged into less target tokens, and every target token is an informative weighted average of the original output tokens. All the target tokens, or the merged tokens, are the final attention output tokens for the LTM-Transformer block, which are fed to an MLP to produce the output of the LTM-Transformer block as illustrated by Figure 1.
@@
Such an idea of informative token merging can also be viewed from the perspective of Information Bottleneck (IB). Let $X$ be the input image. Let $Z$ be the original attention output tokens, which are merged into the merged tokens denoted by $\tilde{X}$, and let $Y$ be the ground truth training labels for a classification task. $\tilde{X}$ has fewer tokens than $Z$. The principle of IB is to maximize the mutual information between $\tilde{X}$ and $Y$ while minimizing the mutual information between $\tilde{X}$ and $X$. That is, IB encourages the network to learn the merged tokens more correlated with the class labels while reducing their correlation with the input. Extensive empirical and theoretical works have evidenced that models respecting the IB principle enjoy compelling generalization. With the informative token merging process in LTM-Transformer, the merged tokens $\tilde{X}$ are the informative aggregation of the original attention output tokens $Z$, so $\tilde{X}$ are less correlated with the training images and in this manner the IB principle is better adhered. This is reflected in Table VI and Table VII in Section 4.4.2, where a model for ablation study with token merging, ToMe, enjoys less IB loss than the vanilla transformer, MobileViT-S. This observation indicates that the IB principle is better respected by the token merging process in ToMe. In order to further decrease the IB loss, we propose an Information Bottleneck (IB) inspired token merging process, where a LTM-Transformer block generates an informative token merging mask, which reduces the IB loss for visual transformers. For example, our model termed “LTM-MobileViT-S” in Table VI and Table VII is the visual transformer with the IB loss reduced by replacing all the transformer blocks in MobileViT-S with LTM-Transformer blocks so that more informative merged tokens are generated by the proposed informative token merging process. While ToMe hurts the prediction accuracy compared to the vanilla model, our LTM-Transformer enjoys even higher top-1 accuracy than the vanilla MobileViT-S, and we have the same observations for MobileViT-XS and EfficientViT.
Contributions. The contributions of this paper are presented as follows.
@@
First, we present a novel and compact transformer block termed Transformer with Information Bottleneck inspired Token Merging, or LTM. Our LTM block generates an informative token merging mask which reduces the IB loss. The LTM blocks can be used to replace all the transformer blocks in many popular vision transformers, rendering compact vision transformers with competitive performance. The effectiveness of LTM is evidenced by replacing all the transformer blocks in popular vision transformers, including MobileViT [18], EfficientViT [6], ViT [3], and Swin [4], with LTM blocks, for image classification, object detection and instance segmentation tasks.
@@
Second, we propose an informative token merging process for vision transformers, which can reduce the IB loss. As a first step, we derive a novel and separable variational upper bound for the IB loss associated with token merging, which is $I(\tilde{X}(G),X)-I(\tilde{X}(G),Y)$ where $I(\cdot,\cdot)$ denotes mutual information and $G$ is the token merging mask in LTM. We then view a transformer with multiple LTM blocks as an iterative process for the reduction of the IB loss by gradient descent, and every LTM block simulates one-step gradient descent on the variational upper bound for the IB loss. Inspired by this understanding, the token merging mask at the current layer is generated from the token merging mask at the previous layer and the input tokens at the current layer by a learnable mask module, following the formula of gradient descent as in (3) in Section 3.2. As a result, such informative token merging process generated in a network with LTM blocks enjoys reduced IB loss, which is evidenced in our ablation study in Section 4.4.2. Due to the separability of the variational upper bound for the IB loss, a neural network with LTM blocks can be trained in an end-to-end manner with standard SGD.
Visualization for Informative Token Merging. Figure 4 illustrates the informative token merging process by LTM with a comparison to LTMP. We remark that LTMP uses uniform merging weights, so their token merging process would not weight informative tokens property, resulting in their inferior performance in classification.
It is worthwhile to mention that our LTM models can be either fine-tuned from pre-trained backbones or trained from scratch. As evidenced in Table I, our LTM models always outperform the current state-of-the-art token merging methods, including the fine-tuning-based method LTMP [14], when fine-tuned for the same number of epochs. We remark that as shown in Table VI, the baseline token merging method, ToMe, and LTMP, can already reduce the IB loss. By replacing all the transformer blocks with our LTM blocks, the networks with LTM exhibit even smaller IB loss and enjoy higher classification accuracy and less FLOPs, either trained from scratch or fine-tuned from pre-trained models. Furthermore, as shown in Table II, our LTM models also outperform all the competing token merging methods when trained from scratch. Importantly, extensive experiment results on various computer vision tasks demonstrate the compelling performance of LTM networks compared to the competing baselines.
@@
This paper is organized as follows. The related works in efficient vision transformers and compression of vision transformers by pruning or token merging are discussed in Section 2. The formulation of LTM is detailed in Section 3. The effectiveness of LTM is demonstrated in Section 4 for image classification, object detection and instance segmentation tasks, by replacing all the transformer blocks of various popular vision transformers, including MobileViT [18], EfficientViT [6], ViT [3], and Swin [4], with LTM blocks. We conclude the paper in Section 5. We use $[n]$ to denote natural numbers between $1$ and $n$ inclusively.
@@ 2.1 Efficient Visual Transformers
Visual transformer models have recently achieved superior performance on a variety of computer vision applications [3, 4, 19, 5, 20]. However, visual transformers often encounter high computational demands due to the quadratic complexity of the point-wise attention and numerous Multi-Layer Perceptron (MLP) layers. To mitigate the challenges of high computational costs, various strategies have been developed [5, 2], primarily aimed at refining the network architectures and incorporating sparse mechanisms for efficient computation. These include the integration of convolutions into transformer networks [18, 6, 21], the use of knowledge distillation for training more efficient transformers [22, 23, 24], and compressing existing visual transformers with methods such as pruning [25, 26, 27]. Techniques for compressing visual transformers generally fall into three categories: (1) Channel Pruning, which targets the elimination of superfluous heads and channels within ViT blocks [25, 28, 29]; (2) Block Pruning, which involves removing redundant transformer blocks [30, 26]; (3) Token Pruning and Token Merging, which prune less important tokens and merge similar tokens in the input of transformer blocks [16, 27, 15, 31].
In this paper, we focus on learning to merge tokens guided by the information bottleneck theory of deep learning and primarily review existing works on Token Pruning and Merging [31, 16, 15, 14, 13]. DynamicViT [16] observes that the prediction in visual transformers is only based on a subset of the most informative tokens and proposes a hierarchical token sparsification framework to prune redundant tokens. ToMe [15] proposes a graph-based matching algorithm that combines similar tokens in each visual transformer block of a pre-trained visual transformer. LTMP [14] learns threshold masking modules that dynamically determine which tokens to merge and prune in a unified framework similar to DynamicViT. ToFu [13] also combines token pruning and token merging. Instead of averagely merging similar tokens, ToFu proposes a conventional average merging module to improve the quality of merged tokens.
@@ 2.2 Related Works about Information Bottleneck
[32] provides the first in-depth analysis of conventional information bottleneck (IB) theories and deep learning to establish the connection between the nonlinearity of neural networks and the compression phase of training. Building on the theory of IB, [33] proposes a probabilistic attention module reducing mutual information between the input and the masked representation while increasing mutual information between the masked representation and the task label. Further exploring the mechanics of IB in deep learning, [34] finds that self-attention mechanisms can be interpreted as iterative steps in optimizing the IB objective, which explains the advantages of self-attention in learning robust representation. Distinct from most existing methods that implicitly incorporate the IB principle, our work adopts a direct and innovative approach. We aim to optimize a novel and separable variational upper bound of the IB loss with a learnable token merging method. The proposed LTM-Transformer leads to compelling performance on many popular visual transformer architectures with lower computation cost, benefiting from the learnable token merging mechanism guided by the IB principle.
(a) LTM block for regular transformers, such as ViT and Swin.
(b) LTM block for efficient transformers, such as MobileViT and EfficientViT.
@@ 3.1 Token Merging by Learnable Token Merging Masks
Given the input feature tokens $X\in\mathbb{R}^{N\times D}$ where $N$ is the number of tokens and $D$ is the token dimension, the LTM block first applies the self-attention module on the input feature tokens by $Z=\text{ATTN}(X)\in\mathbb{R}^{N\times D}$, where $\text{ATTN}(\cdot)$ is the regular QKV self-attention operation [35]. As illustrated in Figure 1, every LTM block has a learnable mask module that generates the token merging mask $G^{(\ell)}$ where $\ell$ is the index of the current layer or block. The LTM block merges the $N$ tokens of $Z$ into $P$ tokens with $P<N$ by multiplying $Z$ with the token merging mask $G^{(\ell)}\in\mathbb{R}^{N\times P}$. We set $P=\lceil r\times N\rceil$, where $r\in(0,1)$ is termed the compression ratio for LTM, and a smaller $r$ renders less merged tokens after token merging. The token merging mask $G^{(\ell)}$ of the $\ell$-th transformer block is generated by the token merging mask $G^{(\ell-1)}$ of the previous layer and the feature tokens $Z$, which is motivated by reducing the IB loss and detailed in Section 3.2. The token merging mask $G^{(1)}$ for the first transformer block is generated by applying an existing learnable token merging method, LTMP [14], which generates a binarized token merging mask $M\in[0,1]^{N\times P}$ using Gumbel-Softmax with $N\times P$ learnable parameters. After obtaining the merging mask $G^{(\ell)}$, the features tokens of $Z$ are merged into $P$ tokens by $\tilde{X}(G^{(\ell)})=\left(Z^{\top}G^{(\ell)}\right)^{\top}\in\mathbb{R}^{P\times D}$, which is then passed to the following MLP layers in the transformer block.
@@
In addition to merging tokens in regular transformer blocks such as ViT [35] and Swin [4], the LTM-Transformer block can also be applied to efficient transformer blocks widely applied in efficient visual transformer architectures such as MobileViT [18] and EfficientViT [6]. Regular transformer blocks obtain the output by sequentially applying the attention operation and the MLP on the input feature tokens. However, efficient transformer blocks usually contain residual connections following the design of residual connections in Convolutional Neural Networks (CNNs). That is, these blocks maintain the same shapes for the input $X$ and the self-attention output $Z$ and concatenate them to produce the output features of the current transformer block. As a result, we cannot only merge the tokens of $Z$. Instead, our LTM-Transformer block merges the tokens of both $X$ and $Z$ so that the number of merged tokens for $X$ and $Z$ is the same. To this end, we apply the same token merging mask $G^{(\ell)}$ to merge both $X$ and $Z$. As a result, the compressed $X$ and $Z$ are of the same shape after the token merging process, and they can still be concatenated, which is illustrated in Figure 1(b). In addition, transformer blocks in the efficient visual transformers are usually accompanied with convolution operations so that they need to maintain the feature tokens in a three-dimensional format $X\in\mathbb{R}^{H\times W\times D}$ as illustrated in Figure 1(b). To apply our token merging method on efficient transformer blocks, we set the number of merged tokens after token merging as $P=H^{\prime}\times W^{\prime}$, where $r$ is the compression ratio, and $H^{\prime}=\lceil H\times\sqrt{r}\rceil,W^{\prime}=\lceil W\times\sqrt{r}\rceil$. Therefore, the merged tokens can still be reshaped into three-dimensional features for later convolution operations. The analysis about the inference computation cost, or the FLOPs, of the LTM transformer block for token merging in both regular transformers and efficient transformers as illustrated in Figure 1 is detailed below.
@@
Computation Cost Analysis of LTM-Transformer for Token Merging. We hereby analyze the additional inference computation cost, or the FLOPs, of the LTM transformer block for token merging in both regular transformers and efficient transformers as illustrated in Figure 1. Let $D$ be the dimension of input tokens and $N$ be the number of tokens. The FLOPs of the token merging in an LTM transformer block in regular visual transformers is $6CDP+3C+ND^{2}+NDP$, where $6CDP+3C+ND^{2}$ is the FLOPs for calculating the merging mask and $NDP$ is the cost for applying the merging mask on the input tokens. In the LTM transformer block of efficient visual transformers, the additional FLOPs of the token merging is $6CDP+3C+ND^{2}+2NDP$, since the merging mask will be applied to both the input tokens and the merged tokens.
@@
3.2 Generating Token Merging Mask by Reducing the Variational Upper Bound for the IB Loss